Define composite functions. For the functions f(x)= x/x+3 and g(x)= 2/x,find f ◦ g
,Also determine the domain for this new function.

Answer by greenestamps(13200)



With the functions you show, f(x) is a constant function:

f(x) = x/x+3 = 1+3 = 4

Then, since f(x) is a constant function, (f ◦ g)(x)=4.

ANSWER: (f ◦ g)(x)=4

I suspect that is not the problem you intended to post. If not, then try re-posting it with proper use of parentheses.